OpenStudy (anonymous):

Find the arc length and area of the circular sector if the radius of the circle is 6 inches and the measure of the central angle is pi/6.

OpenStudy (campbell_st):

arc length formula is \[l = r \theta\] l = arc length, r = radius, theta is the subtended angle in your question \[l = 6 \times \frac{\pi}{6}\] Hope you can finish the answer